Solve the ff. problems?


5. A 3.0 L sample of CO 2 has a pressure of 4.00 atm and a temperature of 30°C. What is the

pressure when the volume of the sample is increased to 10.0 L, and a temperature is

increased to 250°C with no change in moles of gas?


6. A 25-L balloon contains 1.12 moles of gas. How many moles of gas would be required to

expand the balloon to 37 L at constant temperature and pressure?


7. How many moles of H 2 S are contained in a 350 ml bulb at 45°C and 150 kPa?s.




1
Expert's answer
2021-04-08T09:14:19-0400

5)

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}\\\frac{(3)(4)}{273+30}=\frac{10p_2}{273+250}\\p_2=2.07\ atm"

6)


"(1.12)(25)=(37)\nu\\\nu=0.758\ mole"

7)


"pV=\nu RT\\(150000)(0.00035)=(8.31)(25+273)\nu\\\nu=0.0212\ mole"